OpenStudy (anonymous):

Line A passing through the origin also passes through point (2, 3) and is parallel to line B, which passes through the point (7, 6). What is the equation of line B in point-slope form?

OpenStudy (anonymous):

A=y-3=3/2(x-2) B=y-6=3/5(x-7) C=y-6=3/2(x-7) D=y-3=3/5(x-3)

OpenStudy (anonymous):

@aum @Godlovesme

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

What you wanna do is find the slope of the line formed by the first point. That will be the slope for both Line A and B, because they are parallel to each other. Then, use the point-slope formula to fine the equation for Line B.

OpenStudy (anonymous):

I dont know the point slope formula

OpenStudy (anonymous):

y – y1 = m(x – x1)?

OpenStudy (ipwnbunnies):

Yeah, that one lol.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

so 2-7?

OpenStudy (anonymous):

wtf is m=

OpenStudy (ipwnbunnies):

Well, first, you gotta mind the slope of Line A.

OpenStudy (ipwnbunnies):

Using slope formula.

OpenStudy (anonymous):

I need a graph

OpenStudy (anonymous):

I brb this pellet giving me a head ache

OpenStudy (ipwnbunnies):

Lmao, aight. Take your time.

OpenStudy (anonymous):

so slope formula is y^2-y^2/x^2-x^1

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Unanimoose

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

It is C

OpenStudy (ipwnbunnies):

Yeah, use that to find the slope of line A. The two points are (0,0) and (2,3)

OpenStudy (ipwnbunnies):

Hold on.

OpenStudy (ipwnbunnies):

It is C. Good job c:

OpenStudy (anonymous):

yup